# A Fractional Calculus Framework for Open Quantum Dynamics: From Liouville to Lindblad to Memory Kernels

### v1

@@ V Conclusion and Outlook
This unified picture clarifies how fractional differentiation introduces structured memory directly at the generator level, producing CPTP yet non-divisible dynamics that interpolate smoothly between unitary, Markovian, and strongly non-Markovian regimes. It connects the algebraic tails of long-time relaxation with the complete monotonicity of the subordination kernel and explains why fractional dynamics form a structured subclass of memory-kernel quantum master equations.
From an algorithmic perspective, the same structure offers a practical route to simulation. The Mittag–Leffler propagator can be approximated deterministically via polynomial quantum signal processing or stochastically through subordination sampling over operational times. Both approaches realize long-memory quantum dynamics using only Markovian primitives, avoiding explicit time-history storage while preserving complete positivity. This dual deterministic–stochastic formulation establishes a bridge between theoretical open-system analysis and scalable quantum algorithms.
Although the present analysis focused on the exactly solvable pure-dephasing limit at zero temperature, the fractional framework itself imposes no such restriction. Because the subordination construction remains valid for any GKSL generator, the same approach can incorporate amplitude damping, population relaxation, and multiqubit tensor-product generators while preserving complete positivity. Preliminary tests also indicate that the fitted fractional parameters exhibit smooth temperature dependence, suggesting that a single fractional order can continue to describe long-memory behavior for $k_{B}T/\omega_{c}\lesssim 1$.
@@ Appendix C Subordination for the Fractional Master Equation
The subordination representation used in Sec. III.3 follows the classical theory of subordinated semigroups and inverse-stable processes Bochner1955; Feller1971; BaeumerMeerschaert2001; MeerschaertSikorskii2012.
@@
where $\hat{\rho}(s)$ is the Laplace tramsform of $\rho(t)$
@@
Appendix D Optimization of fractional parameters
@@
which is compared to the numerically exact Nakajima–Zwanzig coherence $u_{\mathrm{NZ}}(t)=\exp[-Q(t)]$ obtained from the structured spectral density $J(\omega)$. The parameters $(\alpha,\lambda)$ are determined by a least–squares fit over a finite time window $t_{\mathrm{start}},t_{\mathrm{end}}]$:
@@
The fitting window is guided by the bath correlation time. In our dimensionless units we set $\omega_{c}=1$, so the bath correlation time is of order $\tau_{B}\sim 1/\omega_{c}=1$. More quantitatively, we estimate $\tau_{B}$ from the decay of the bath correlation function $C(t)=\langle B(t)B(0)\rangle$ (or equivalently from $Q^{\prime\prime}(t)$) as the time at which $|C(t)|$ has dropped below $e^{-1}|C(0)|$. The fitting windows are then chosen as $t_{\mathrm{start}}\approx 2\tau_{B}$ and $t_{\mathrm{end}}\in[20\tau_{B},60\tau_{B}]$, which in practice corresponds to $2/\omega_{c},20/\omega_{c}][$$2/\omega_{c},60/\omega_{c}]$ in Fig. 4. This ensures that the optimized $(\alpha,\lambda)$ are extracted from a regime where the bath has largely decorrelated but before finite-cutoff recurrences become important.

### v2

@@ IV Conclusion and Outlook
This unified picture clarifies how fractional differentiation introduces structured memory directly at the generator level, producing CPTP yet non-divisible dynamics that interpolate smoothly between unitary, Markovian, and strongly non-Markovian regimes. It connects the algebraic tails of long-time relaxation with the complete monotonicity of the subordination kernel and explains why fractional dynamics form a structured subclass of memory-kernel quantum master equations. Moreover, by linking the fractional resolvent to NZ kernels, HEOM self-energies, and influence-functional descriptions, the present framework situates fractional dynamics as a coarse-grained, analytically transparent complement to existing microscopic approaches rather than a replacement for them.
From an algorithmic perspective, the same structure offers a practical route to simulation. The Mittag–Leffler propagator can be approximated deterministically via polynomial quantum signal processing or stochastically through subordination sampling over operational times. Both approaches realize long-memory quantum dynamics using only Markovian primitives, avoiding explicit time-history storage while preserving complete positivity. This dual deterministic–stochastic formulation establishes a bridge between theoretical open-system analysis and scalable quantum algorithms. Such reduced, generator-level models may be particularly valuable when fully microscopic solvers—such as HEOM, QUAPI, or tensor-network path-integral methods—become computationally expensive, yet one still seeks controlled long-time behavior or interpretable memory parameters.
Although the present analysis focused on the exactly solvable pure-dephasing limit at zero temperature, the fractional framework itself imposes no such restriction. Because the subordination construction remains valid for any GKSL generator, the same approach can incorporate amplitude damping, population relaxation, and multiqubit tensor-product generators while preserving complete positivity, thus apply to models with dissipation, pumping, and non-commuting system–bath couplings, as clarified in Sec. II.2.4. The pure-dephasing model was chosen solely because it provides an exact analytical reference against which the fractional formulation can be benchmarked without numerical ambiguity. A natural extension is to apply this framework to dissipative models (with energy relaxation) where only numerical references such as HEOM or QUAPI are available, which will be explored in future work.
@@ Appendix C Subordination for the Fractional Master Equation
The subordination representation used in Sec. II.2.3 follows the classical theory of subordinated semigroups and inverse-stable processes Bochner1955; Feller1971; BaeumerMeerschaert2001; MeerschaertSikorskii2012.
@@
where $\hat{\rho}(s)$ is the Laplace transform of $\rho(t)$
@@
Appendix D Predictive Determination of $(\alpha,\lambda)$ from the Spectral Density
This appendix provides the derivations underlying the predictive rules for $(\alpha_{\rm pred},\lambda_{\rm pred})$ summarized in Table 2. Our goal is to extract the fractional order $\alpha$ and amplitude $\lambda$ directly from microscopic inputs, which are the short- and long-time asymptotic structures of the dephasing kernel $Q(t)$ induced by the spectral density $J(\omega)$.
We begin with the exact coherence $u(t)=e^{-Q(t)}$ of the pure–dephasing spin–boson model and compare it to its fractional analogue
$u_{\alpha}(t)=E_{\alpha}(-\lambda t^{\alpha}),$ (72)
where $E_{\alpha}$ denotes the Mittag–Leffler function. A crucial feature of $E_{\alpha}$ is that it exhibits two distinct asymptotic regimes, which must be used appropriately depending on the bath structure:
  • (i)
    Small-argument / moderate-time expansion ($\lambda t^{\alpha}\ll 1$):
    $E_{\alpha}(-\lambda t^{\alpha})$ $=1-\frac{\lambda}{\Gamma(1+\alpha)}t^{\alpha}+\mathcal{O}(t^{2\alpha})$
    $\approx\exp[-\frac{\lambda}{\Gamma(1+\alpha)}t^{\alpha}\Big].$ (73)
    This behavior mimics a stretched exponential and is the form relevant for matching sub–Ohmic baths, whose exact coherence also decays as a stretched exponential.
  • (ii)
    Large-argument / long-time asymptotic ($\lambda t^{\alpha}\gg 1$):
    $E_{\alpha}(-\lambda t^{\alpha})\sim\frac{1}{\lambda\,\Gamma(1-\alpha)}\,t^{-\alpha}~~~~(t\rightarrow\infty).$ (74)
    This algebraic decay is the correct asymptotic structure for matching Ohmic baths (which yield power-law decoherence) and super–Ohmic baths (where the coherence relaxes to its plateau with a power-law correction).
The correct identification of $(\alpha,\lambda)$ therefore depends on whether the exact coherence $u(t)$ (determined by the long-time form of $Q(t)$) is stretched exponential (sub–Ohmic) or algebraic (Ohmic or super–Ohmic). In the sections below, we match the appropriate Mittag–Leffler regime to the asymptotic expansions of $Q(t)$ that follow from the low-frequency structure of the spectral density shown in Eq. (31).
Universal short-time regime
The short-time expansion of $Q(t)$ is universal for all bath types. Expanding the cosine in Eq. (30) gives
$Q(t)=A_{\chi}\,t^{2}+\mathcal{O}(t^{4}),~~\text{with}~~A_{\chi}=\frac{\eta\,\Gamma(\chi+1)\,\omega_{c}^{2}}{\pi}.$ (75)
Thus the exact coherence behaves as
$u(t)$ $=\exp(-A_{\chi}\,t^{2}+\mathcal{O}(t^{4}))$
$=1-A_{\chi}t^{2}+\mathcal{O}(t^{4}).$ (76)
The Mittag–Leffler expansion, Eq. (73), matches the curvature only when
$\alpha_{\rm pred}=2,~~\lambda_{\rm pred}=2A_{\chi}~.$ (77)
These relations provide the analytic short-time assignments used in Figure 6 and in the main text. They require only the microscopic coefficient $A_{\chi}$, and involve no fitting.
Sub–Ohmic baths ($0<\chi<1$)
For $J(\omega)\propto\omega^{\chi}$ with $0<\chi<1$, the long-time dephasing is dominated by low-frequency modes. The asymptotic form is
$Q(t)\sim C_{\chi}\,t^{1-\chi}~~~~(t\to\infty),$ (78)
with $C_{\chi}$ determined by the infrared part of the spectral density. The corresponding coherence decays as
$u(t)$ $\sim e^{-C_{\chi}t^{1-\chi}}$
$\sim 1-C_{\chi}t^{1-\chi}+\mathcal{O}(t^{2-2\chi}).$ (79)
Matching the leading power-law behavior of $E_{\alpha}(-\lambda t^{\alpha})$ in Eq. (73) yields
$\alpha_{\rm pred}=1-\chi,~~\lambda_{\rm pred}=C_{\chi}\,\Gamma(2-\chi)~.$ (80)
Ohmic bath ($\chi=1$)
For the Ohmic case,
$J(\omega)=\eta\,\omega\,e^{-\omega/\omega_{c}},$ (81)
the exact long-time form is
$Q(t)\sim\frac{2\eta}{\pi}\,\ln t+\text{const}~~\rightarrow~~u(t)\sim t^{-2\eta/\pi}.$ (82)
Matching the Mittag–Leffler asymptotic in Eq. (74) gives
$\alpha_{\rm pred}=\frac{2\eta}{\pi},~~\lambda_{\rm pred}=\frac{1}{\Gamma(1-\alpha_{\rm pred})}~.$ (83)
Super–Ohmic baths ($\chi>1$)
For $\chi>1$, the dephasing saturates at a finite value:
$Q_{\infty}=\frac{2\eta}{\pi}\,\Gamma(\chi-1),~~u_{\infty}=e^{-Q_{\infty}}.$ (84)
The asymptotic correction takes the form
$Q(t)=Q_{\infty}-D_{\chi}\,t^{-(\chi-1)}+\cdots,$ (85)
so that
$u(t)-u_{\infty}\sim u_{\infty}\,D_{\chi}\,t^{-(\chi-1)}.$ (86)
Define the plateau-normalized coherence $v(t)$ as Eq. (45), then
$v(t)\sim\frac{u_{\infty}}{1-u_{\infty}}\,D_{\chi}\,t^{-(\chi-1)}.$ (87)
Matching the Mittag–Leffler form Eq. (74) to above $v(t)$ expression yields
$\alpha_{\rm pred}=\chi-1,~~\lambda_{\rm pred}=\frac{e^{Q_{\infty}}-1}{D_{\chi}\,\Gamma(2-\chi)}.$ (88)
Summary
The microscopic short- and long-time structures of $Q(t)$ fully determine the fractional parameters $(\alpha_{\rm pred},\lambda_{\rm pred})$. Short-time matching always gives $(\alpha,\lambda)=(2,2A_{\chi})$, while the long-time exponents and amplitudes follow directly from the infrared behavior of the bath through the power $\chi$. The results for all three regimes match those listed in Table 2, and are precisely the values used in the parameter-free comparisons of Sec. III.B and Figure 6.
Appendix E Optimization of fractional parameters
@@
which is compared to the numerically exact NZ coherence $u_{\mathrm{NZ}}(t)=\exp[-Q(t)]$ obtained from the structured spectral density $J(\omega)$. The parameters $(\alpha,\lambda)$ are determined by a least–squares fit over a finite time window $t_{\mathrm{start}},t_{\mathrm{end}}]$:
@@
The fitting window is guided by the bath correlation time. In our dimensionless units we set $\omega_{c}=1$, so the bath correlation time is of order $\tau_{B}\sim 1/\omega_{c}=1$. More quantitatively, we estimate $\tau_{B}$ from the decay of the bath correlation function $C(t)=\langle B(t)B(0)\rangle$ (or equivalently from $Q^{\prime\prime}(t)$) as the time at which $|C(t)|$ has dropped below $e^{-1}|C(0)|$. The fitting windows are then chosen as $t_{\mathrm{start}}\approx 2\tau_{B}$ and $t_{\mathrm{end}}\in[20\tau_{B},60\tau_{B}]$, which in practice corresponds to $2/\omega_{c},20/\omega_{c}][$$2/\omega_{c},60/\omega_{c}]$ in Figure 7. This ensures that the optimized $(\alpha,\lambda)$ are extracted from a regime where the bath has largely decorrelated but before finite-cutoff recurrences become important.